OpenStudy (anonymous):

did I do this right? 1/2 - 7/8 = -3/8

OpenStudy (anonymous):

the / is division sign, i also would like to know how to solve 2y^5(10y^2+4y)

OpenStudy (anonymous):

solve? you haven't written an equation.

OpenStudy (anonymous):

yes my equation is 2y^5(10y^2+4y)

OpenStudy (anonymous):

@pgpilot326

OpenStudy (anonymous):

how would i solve this

OpenStudy (skullpatrol):

Where is the equals sign?

OpenStudy (skullpatrol):

An equation is a statement formed by placing an equals sign between two numerical or variable expressions.

OpenStudy (anonymous):

that's the way it is srry

OpenStudy (anonymous):

I think you would use the distributive property right

OpenStudy (skullpatrol):

yes, that would simplify it

OpenStudy (skullpatrol):

but not SOLVE it

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so how would i simplify it step by step ?

OpenStudy (skullpatrol):

$$\Huge a(b+c) = ab+ac$$ $$\Huge 2y^5(10y^2 +4y)$$ $$\text{Let a =}2y^5$$

OpenStudy (anonymous):

2y5(b+c)=ab+ac

OpenStudy (skullpatrol):